OpenStudy (anonymous):

1. local max, local min, and saddle point of f (x, y) = x^3 + 8y^3 - 6xy 2. A rectangular box, without a lid made 12m^2 of cardboard. Find the max volume of such a box?

OpenStudy (anonymous):

Partial differentiation wrt x \[\delta f / \delta x = 6x ^{2}-6y \] Equating this to 0 we get => x^2 = y Partial differentiation wrt y \[\delta f / \delta y = 48x ^{2}-6y \] Equating this to 0 we get => 4y^2 = x Solving this two equations we get (0,0), (0.63, 0.4) and (0.63, -0.4) Now \[\delta ^{2} f / \delta x ^{2} = 12x\] And \[\delta ^{2} f / \delta y ^{2} = 48y\] And \[\delta ^{2} f / \delta y \delta x = -6\] To discover the nature of the turning point @ point (x,y), the nature of the roots of the following equations are to be checked for lamda \[(\delta ^{2} f / \delta x ^{2} - \lambda)(\delta ^{2} f / \delta y ^{2} - \lambda) - (\delta ^{2} f / \delta y \delta x)^2 = 0\] - at (0,0) \[(-\lambda)((-\lambda) - (-6)^{2} = 0 => \lambda = \pm6\] Since, the roots are of opposite sign hence it is a saddle point at (0.63,0.4) \[(12x-\lambda)((48y-\lambda) - (-6)^{2} = 0 \] Both of these roots are positive, hence there is local minimum, Similarly for ((0.63,-0.4), both roots are negative, hence there is a local maximum

OpenStudy (anonymous):

For question 2, let the three sides of the box be x, y and z. Hence the volume is f(x,y,z) = xyz The constraint function The cupboard side is 12m^2. Now for all side the total area of the walls is g(x,y,z) = (2xy + 2yz + 2zx - 12) Hence if the Lagrangian function be L and multiplier be lambda, then \[L = f(x,y,z) - \lambda g(x,y,z)\] The partial differentiations are to be equated to zeroes - \[\delta L / \delta x = yz - 2\lambda (y+z) = 0 => \lambda = yz / 2(y+z)\] Similarly, \[\delta L / \delta y = zx - 2\lambda (z+x) = 0 => \lambda = zx / 2(z+x)\] \[\delta L / \delta z = xy - 2\lambda (x+y) = 0 => \lambda = xy / 2(x+y)\] Solving all these reaches => x=y=z Again, \[\delta L / \delta \lambda = 2xy + 2yz + 2zx - 12 = 0 => 6x^2 = 12 => x = \sqrt{2}\] Hence\[V = (\sqrt{2})^{3} = 2\sqrt{2} \approx 2.83m\]

OpenStudy (anonymous):

sorry the unit should be m^3

OpenStudy (anonymous):

Please let me know if you need any explanation of any step

OpenStudy (anonymous):

owh i'll need time to read up 1st. because i was thrown with this questions by my lecturer when he haven even started the calculus topic. anyway thank you very much. =)

OpenStudy (anonymous):

actually these are little advanced topics, calculus generally should start with idea of limit

OpenStudy (anonymous):

erm, owh then i should rephrase it as, he started the basics of calculus haha. still at early stages. probably he wanted us to do some self study.

OpenStudy (anonymous):

yeah, cheers...

OpenStudy (anonymous):

just one thing can you see the posted equation in the correct form

OpenStudy (anonymous):

yeah they are very neat. thank you very much for spending your time on it.

OpenStudy (anonymous):

welcome, i think there is some problem with my machine though :)

OpenStudy (anonymous):

erm, based on © 2007 Paul Dawkins' notes, Q1 should give (0,0) as saddle, and (1,1/2) as minimum, while Q2 gives volume of \[3/2\sqrt{6}\]

OpenStudy (anonymous):

I agree with the saddle point before, the note you are referring is correct on local minima. (I made a mistake in taking the partial differentiation wrt x, thank you for pointing out) And I feel the volume is is correct as per my calculation. Can you please give the length of each side of the box as per the note

OpenStudy (anonymous):

erm the question is not from the notes, i was just searching the net for some notes on lagrange multiplier, and i found paul's notes and apply it to the question. i think the difference is just at you're using 2xy+2yz+2xz as the g(x), where the question i mentioned is concerning on a box without lid. sorry for not bolding it. lol

OpenStudy (anonymous):

Hey, you are correct, I missed it. again!!! I jumped quite a few guns here :)